OpenStudy (anonymous):

Y=-4X+5 y=-x-1 write the solutions as an ordered pair

OpenStudy (anonymous):

subtract these equations!

OpenStudy (anonymous):

u can substitue (-x-1)=-4x+5

OpenStudy (anonymous):

0=-4x+5+x+1 0=-3x+6 -3x=-6 x=2

OpenStudy (anonymous):

now u can use the value of x to find y

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

so (2,-3)

OpenStudy (anonymous):

then write it in form (x,y) (2,?) which is ordered pair

OpenStudy (anonymous):

i can't tell u the whole answer

OpenStudy (anonymous):

I GOT IT THANK